Bowser Products operates a small plant in New Mexico that produces dog food in batches of 1,500 pounds. The product sells for $6 per pound. Standard costs for 2018 are:


Standard direct labor cost = $15 per hour

Standard direct labor hours per batch = 10 hours

Standard price of material A = $0.35 per pound

Standard pounds of material A per batch = 800 pounds

Standard price of material B = $0.55 per pound

Standard pounds of material B per batch = 250 pounds

Fixed overhead cost per batch = $500

At the start of 2018, the company estimated monthly production and sales of 50 batches. The company estimated that all overhead costs were fixed and amounted to $25,000 per month. During the month of June 2018 (typically a somewhat slow month), 42 batches were produced (not an unusual level of production for June). The following costs were incurred:

Direct labor costs were $7,800 for 460 hours.

37,500 pounds of material A costing $8,500 were purchased and used.

12,000 pounds of material B costing $5,600 were purchased and used.

Fixed overhead of $23,000 was incurred.

a. Calculate variances for material, labor, and overhead.

b. Prepare a summary of the variances. Does the unfavorable overhead volume variance suggest
that overhead costs are out of control?
Answer #1

a.

Labor variance:

Standard labor cost on monthly basis = Standard Direct labor cost $ per hour * Standard Direct labor hours per batch * estimated number of batches

= $15 * 10 * 50 = $7,500

Jun 2018 actual direct labor costs were 7,800

Hence, Labor cost variance = Standard labor cost - Actual labor cost = 7,500 - 7,800 = - $300

Material variance:

Standard cost of material A= Standard price of material A * Standard pounds of material A per batch * estimated number of batches

= $0.35 per pound * 800 pounds per batch * 50 batches

= $14,000

(800*50 = 40000 pounds)

Standard cost of material B= Standard price of material B * Standard pounds of material B per batch * estimated number of batches

= $0.55 * 250 * 50

= $ 6,875

(250*50 = 12500 pounds)

Hence, material variance for A = Standard material cost of A - Actual cost of A = $14,000 - $8,500 = $5,500

Material variance for B = Standard material cost of B - Actual cost of B = $6,875 - $5,600 = $1,275

Overhead variance:

Standard Fixed Overhead cost = Fixed overhead cost per batch * estimated number of batches

= $500 * 50 = $25,000

Actual Fixed overhead cost = $23,000

Overhead variance = Standard overhead cost - Actual overhead cost = $25,000 - $23,000 = $2,000

b.

Summary of variances

  • Labor cost variance is - $300
  • Material variance for A is $5,500
  • Material variance for B is $1,275
  • Total material variance (A + B) is 6,775
  • Overhead variance is $2,000

Overhead cost is fixed for each batch. Standard overhead cost for each batch is $500. In actual fixed overhead cost of $23,000 is incurred for 42 batches i.e. $547.62 per batch. Hence, unfavorable overhead volume variance exist and this suggest overhead cost are out of standard and out of control.

If you liked my answer, please give like/ thumbs up. Thank you!